OpenStudy (anonymous):

Simplify Please: x^3+8x^2+20x+13/(x+3)(x-2)

OpenStudy (anonymous):

It didn't simplify it completely :/

OpenStudy (anonymous):

It's all good

OpenStudy (jhannybean):

Did you figure it out though, @itsallycatt?

OpenStudy (anonymous):

No I didn't...

OpenStudy (jhannybean):

So your problem follows the "sum of cubes" formula, in that \[(a^3+b^3)= (a+b)(a^2-ab+b^2)\]

OpenStudy (jhannybean):

@jim_thompson5910 over here

OpenStudy (anonymous):

Are you sure that is the equation? I don't see any way in which it can be simplified substantially. Then again I might just be missing something.

OpenStudy (jhannybean):

ohhhh i figured it out, i think haha

OpenStudy (anonymous):

That's what I thought... The answer choices are 1/(x+3)(x+4) 1/(x+3)(x-2) 1/(x+4)(x-2) 1/(x+3)(x+1)

OpenStudy (jhannybean):

multiply (x+3)(x-2) to get a quadratic equation and then use long division to simplify

jimthompson5910 (jim_thompson5910):

I'm assuming all of x^3+8x^2+20x+13 is over (x+3)(x-2) If so, then you can expand out (x+3)(x-2) to get x^2+x-6 then you can use polynomial long division to get the quotient for (x^3+8x^2+20x+13)/(x^2 + x - 6)

OpenStudy (jhannybean):

xD yay jim! haha

OpenStudy (jack1):

(x+3)(x-2) = x^2 + x - 6 and x^3+8x^2+20x+13/ x^2 + x - 6 = (x+7) (x^2 + x - 6) with a remainder of (19x +55)

jimthompson5910 (jim_thompson5910):

something seems missing though because none of the answer choices will match the quotient or the remainder so is there more to the problem?

OpenStudy (anonymous):

I have no ideaa, that was the problem I was having...

OpenStudy (jhannybean):

@Jack1 how did you get your (x+7)(x^2 +x - 6)....

OpenStudy (jack1):

none of these match up... do they have brackets around any of the initial question on the top line? 1/(x+3)(x+4) 1/(x+3)(x-2) 1/(x+4)(x-2) 1/(x+3)(x+1)

OpenStudy (jack1):

@Jhannybean long division

OpenStudy (anonymous):

No they don't...

OpenStudy (jhannybean):

Ohhh. nevermind. Gotcha.

OpenStudy (jack1):

(x^2 + x - 6) times x = x^3 +x^2 -6x goes into x^3+8x^2+20x ..cool nvmd

OpenStudy (jhannybean):

xD yeah this problem is kicking my booty.

OpenStudy (jack1):

@Jhannybean dont feel bad, its not the answerer, its the question, the fact that there's a remainder means that it doesn't simplify into one of the optional answers, so something in the q's wrong, maybe the people who wrote it put a sign wrong... dunno